An inheritance of $40,000 is invested in two municipal bonds which pay 6% and 7% simple annual interest. If the annual interest from both bonds is 2,550 dollars how much is invested in each bond?
Answer by rfer(16322)

.06x+.07(40000-x)=2550
.06x+2800-.07x=2500
-.01x=-300
x=30000 at 6%
40000-x=10000 at 7%